### v2

# Timeseria: an object-oriented time series processing library

Abstract
Timeseria is an object-oriented time series processing library implemented in Python, which aims at making it easier to manipulate time series data and to build statistical and machine learning models on top of it. Unlike common data analysis frameworks, it builds up from well defined and reusable logical units (objects), which can be easily combined together in order to ensure a high level of consistency. Thanks to this approach, Timeseria can address by design several non-trivial issues often underestimated, such as handling data losses, non-uniform sampling rates, differences between aggregated data and punctual observations, time zones, daylight saving times, and more. Timeseria comes with a comprehensive set of base data structures, common data manipulation operations, and extensible models for data reconstruction, forecasting and anomaly detection. It also integrates a powerful plotting engine capable of handling even millions of data points.
keywords:
Python , Time series, Data structures, Forecasting, Reconstruction, Anomaly detection
journal: SoftwareX
Metadata
Current code version v2.0.0
Permanent link to code repository for this code version https://github.com/sarusso/Timeseria/tree/v2.0.0
Permanent link to reproducible capsule https://codeocean.com/capsule/4941409
Code license Apache Licence v2
Code versioning system Git
Code languages Python, JavaScript
Operating environments & dependencies Required environment: Python >= 3.6 Optional environments: Jupyter Lab >=3, Jupyter Notebook >=5 Required Python dependencies: keras >=2.1.3, <3.0.0, matplotlib >=2.1.2, <4.0.0, numpy >=1.19.5, <2.0.0, scikit-learn >=0.2.2, <2.0.0, pandas >=0.23.4, <2.0.0, chardet >=3.0.4, <4.0.0, convertdate >=2.1.2, <3.0.0, lunarcalendar >=0.0.9, <1.0.0, cython >=0.29.17, <1.0.0, requests >=2.20.0, <3.0.0, h5py >=2.10.0, <4.0.0, scipy >=1.5.4, <2.0.0, pyppeteer>=0.2.6, <1.0.0, fitter>=1.7.0, <2.0.0, propertime>=1.0.0, <2.0.0 Optional Python dependencies: tensorflow >=1.15.2, <3.0.0, prophet: >=1.1.1, <2.0.0, pmdarima >=1.8, <2.0.0, statsmodels >=0.12.1, <1.0.0 Optional system dependencies: Chrome or Chromium >= 57 for image-based plots (automatically downloaded if not present)
Documentation https://timeseria.readthedocs.io/en/v2.0.0
Support for questions https://github.com/sarusso/Timeseria/issues
Table 1: Code metadata
1 Motivation and significance
Time series are central to several fields across many disciplines, such as engineering, meteorology, medicine, environmental science, finance, economics, and more. Time series represent the evolution of a phenomena over time, and their analysis is essential to capture the dynamics of the phenomena being studied, understand cause-and-effect relationships, and make predictions.
However, a typical time series processing pipeline — loading a data set, cleaning and plotting it, performing some operations, applying some models and inspecting the results — still feels unnecessarily cumbersome. Scientists, engineers, analysts, and many other professionals spend a considerable amount of time on repetitive tasks and on getting their code to work, instead of focusing on their core analysis.
Current numerical and data analysis frameworks[1][2][3][4] indeed fall short in handling everyday practical challenges when it come to time series data, and perhaps most importantly, in real-world scenarios.
In a recent review[5] that surveyed 40 time series processing packages, within the 17 listed as capable of handling missing data we found that none of them provided a robust solution, but more of a set of workarounds. Similarly, among the packages listed as capable of performing modeling tasks, checking for data consistency was always left to the user.
We also noted a generalized lack of definitions about what constitutes time series data, and how to represent it[6]. This is partly because each format has it own advantages and disadvantages, and partly because the complexity of such data is often underestimated.
Simple vector or matrix data structures are usually the default choice, until discovering at later stages that are just not capable of taking into account most of the typical time series characteristics: data losses are represented ambiguously, calendar arithmetic and daylight saving times are handled poorly, there is no support for variable sampling rates and no clear distinction between punctual observations and aggregated data.
In this context, the usual solution is a series of patches stuck together which eventually backfires, and very few attempts have been made to get the fundamentals right. The TSflex[7] package, which focuses on time-series preeprocessing and feature extraction, is one of such rare examples.
To the best of our knowledge, as of today there is neither research work nor software packages aimed at addressing such issues at their roots.
In this work we presents a novel, object-oriented approach for representing time series data, which we believe can address most of the issues outlined above, together with the software library implementing it: Timeseria.
2 Software description
Timeseria presents itself as a Python library that can be imported and used in interactive data analysis environments (as Jupyter), batch data analysis pipelines, and more structured projects. It supports both univariate and multivariate time series, and provides functionalities for loading, storing, manipulating and plotting data, as well as fitting, evaluating and applying models.
2.1 Software architecture
Figure 1: Timeseria base classes structure
The main modules of Timeseria are the Datastructures, the Operations, the Transformations and the Models, which are schematized in Figure 1 together with their main classes, and described more in detail in the following.
2.1.1 Data structures
The Datastructures module provides all the base data structures: Points, Slots and Series, together with their specialisations.
A Point is just a point in any n-dimensional space, which supports common mathematical operations. A TimePoint is a point in the time dimension, measured in epoch seconds, where the zero corresponds to the midnight of 1st January 1970, UTC. Sub-second precision can be achieved with decimals. A DataPoint is an extension of the point concept, with some data attached. Such data can be of any kind, however Timeseria works best with indexed or key-value (labeled) data. To define data points over time, the DataPoint and TimePoint classes are merged, creating the DataTimePoint.
Since points are intended to represent punctual observations, to represent aggregated data Timeseria defines a new data structure: the Slots. These are “containers” that span from a start to an end (which are again point objects), and that are naturally associated with a unit, which represents their span. Similarly to the points, also for the slots Timeseria defines the DataSlots, TimeSlots and the DataTimeSlots.
Both data points and data slots support a data_loss attribute, which is a key feature in Timeseria. This represents the data loss occurred when computing a new point in resampling (i.e. because its nearest neighbours were far), or when computing a new slot in aggregating (i.e. because of missing punctual observations).
The data loss is a particular case of a more generic concept: the data indexes. These are indicators in the 0 - 1 range of some property of the data attached to a data point or slot. Example data indexes include, besides the data loss, the data reconstructed index, the forecast index, and the anomaly index.
Series are defined as a list of items coming one after another, where every item is guaranteed to be of the same type, and following an order or a succession. Ordering is checked with the standard comparison operators, while it is up to the series items to define a succession logic, if any. In Timeseria, points only implement an ordering, while slots also define a succession: in other words, it is enforced that if a time series is slotted, then it must be dense, without any hole.
Series can carry virtually any data type, but Timeseria focuses on the time domain and thus defines a dedicated specialization: the TimeSeries. This is a collection, again in order or succession, of points or slots over time (e.g. DataTimePoints or DataTimeSlots). Time series also have a resolution attribute, which represents the temporal resolution. This concept is similar to a sampling interval, but generalised in order to make it compatible with slots as well. For example, a point time series sampled at 1-minute intervals has resolution corresponding to the time unit “60s”, while a one-day slots time series has resolution corresponding to the time unit “1D”. Time series with a variable sampling rate (e.g. due to a non-constant sampling process) report instead “variable” as their resolution.
Lastly, one of the main architectural features of Timeseria is to allow time units to have variable length. This allows to effectively model months with a variable number of days, and days with a variable number of hours (due to the daylight saving time). It does indeed not hold true that a time unit of 24 hours (“24h”) is equal to a time unit of one day (“1D”) : the first is always exactly 24 hours, while the second can assume different lengths: 24, 24 or 25 hours, depending whether there is an ongoing change in the daylight saving time, and in which direction.
2.1.2 Transformations
Transformations change the temporal resolution of a time series, and modify its data accordingly. The base Transformation class provides a generic transformation concept, which is then specialized in the Resampler and the Aggregator.
The first transforms a point series in another point series with a different sampling interval, while the second transforms a point series in a slot series, aggregating its data in slots of a given unit. The main difference between the two is that while a resampler is supposed to produce an alternative “view” of the original data, an aggregation is instead intended to compute statistical indicators, the default one being the average.
Transformations also detect missing data, impute it using an Interpolator while setting the data_loss attribute accordingly, and recompute other pre-existent data indexes in order to bring them forward.
2.1.3 Operations
Operations are meant to manipulate the data of a series, and can return other series, a scalar, a list of items, or any other valid data type. An operation, if returning a series, it never changes its temporal resolution, which is a task reserved for the transformations. Timeseria provides several built-in operations, as computing the average, minimum and maximum values, merging two series, applying an offset, normalizing the data, and more. Custom operations can be defined as well.
2.1.4 Models
Timeseria defines a base Model class which can represent different kind of models, and in particular parametric models, where there is a (potentially very large) number of parameters that can be either set manually or learned (fitted) from the data.
All models expose a predict(), apply() and evaluate() methods. Parametric models also provide a save() method to store the parameters of the model, and an optional fit() method if the parameters are to be learned form the data. In this case also a cross_validate() method is available.
Models also enforce resolution and data consistency within the various steps, and are sub-dived in three main categories: forecasters, reconstructors, and anomaly detectors.
  • 1.
    Forecasters make predictions for n steps ahead, either using their internal logic to predict all steps in one go, or by recursively re-applying the same logic n times on previously predicted data. Forecasters can operate with or without an input window. Edge cases (as applying a forecaster on a time series without enough data if using a window) are automatically handled by the base Forecaster class, which also adds the forecast data index.
  • 2.
    Reconstructors work on reconstructing missing data, or in other words to fill gaps. Gaps need a “next” element to be defined, and are detected using the data_loss attribute. By default, only gaps with a full data loss are reconstructed. Reconstructors can operate with or without an input window, on either sides of the gap. Edge cases (as applying a reconstructor on a time series without enough data before or after a gap) are automatically handled by the base Reconstructor class, which also adds the reconstructed data index.
  • 3.
    Anomaly detectors aim at spotting anomalies in the data. There are several ways to achieve this goal, and a structured exposition of the topic is out of scope for this article. As of today, Timeseria implements the so called model-based anomaly detection[8] only. In this mode, anomalies are detected by evaluating the error between the actual values and the values predicted by a model: if this is “too big”, then it is considered as a symptom of anomaly. Timeseria provides the ModelBasedAnomalyDetector class for such task, which implements all of the model-based anomaly detection logic. Any model sub-classing the base Model class can be plugged-in, and in particular all of the Forecasters and Reconstructors, including user-defined ones. Each data point or slot is then assigned with an an anomaly index (as the anomaly data index), representing the likelihood of that point or slot of being anomalous. Such index is computed taking into account the overall error distribution of the model together with some probabilistic considerations (see the documentation for more details).
2.2 Software functionalities
As already introduced, Timeseria provides functionalities for loading, storing, manipulating and plotting data, as well as fitting, evaluating and applying models. In the following, we provide a breakdown of the main ones, together with their respective methods or classes.
2.2.1 Loading and storing data
Timeseria supports loading and storing data directly from the TimeSeries class. The from_csv() method allows to load CSV files with automatic encoding and format detection, and supports many parameters should the auto-detection fail. Field and newline separators, timestamp formats, timestamp columns and which data columns to read can all be set, as well as many others. Time series can be saved either as standard CSV files using the to_csv() method, or as dedicated format with the save() method, that includes some metadata and that fully supports data indexes. This is particularly useful for loading time series within Timeseria itself later on, which can be done using the load() method. Lastly, Pandas DataFrames are supported as well, using the to_df() and from_df() methods.
2.2.2 Manipulating and inspecting data
Data can be manipulated using the transformations and the operations, which can be accessed directly as methods of the TimeSeries class. Besides the two main resample() and aggregate() transformations, Timeseria provides a number of built-in operations which include: min(), max(), avg(), sum(), derivative(), integral(), diff(), csum(), mavg(), normalize(), offset() and rescale() to operate on the time series values, plus filter() to filter by label, slice() to cut a portion of a time series and merge() to merge two or more time series.
The TimeSeries class also supports the square brackets notation for quickly accessing its elements either by position or timestamp (e.g. timeseries[2], timeseries[dt(2015,10,25,6,19,0)]), or for filtering on a specific data label if using key-value data (e.g. timeseries[’temperature’]). Slicing is also possible, either by position or timestamp, using the standard Python slicing notation (e.g timeseries[dt(2015,1,1,0,0,0):dt(2016,1,1,0,0,0)]).
Lastly, plotting a time series can be achieved by simply invoking the plot() method of the TimeSeries class, which generates interactive plots using a JavaScript-based engine. Such plots fully support Timeseria data structures, including the data indexes, and can be saved both as images and as self-contained HTML pages in order to be easily shared. Above a given amount of data, the plotting engine automatically performs an aggregation step, in order to speed up the plotting process and to prevent overwhelming interactive analysis environments (and web browsers) with too much data. In this case, the plot is rendered as a line chart overlying an area chart which represents the dispersion of the original data, thus allowing to retain the visual information about peaks that would be otherwise smoothed.
2.2.3 Fitting, evaluating and applying models
Timeseria provides a set of built-in models to be intended more for demonstrative and/or simple tasks rather than for real applications. These include the PeriodicAverageForecaster, the PeriodicAverageReconstructor and the PeriodicAverageAnomalyDetector, which are all based on computing periodic averages over historical data; the AARIMAForecaster and the AARIMAReconstructor, based on automatic autoregressive integrated moving averages; and the LSTMForecaster and the LSTMAnomalyDetector, which make use of an LSTM neural network.
The true goal of Timeseria is instead to provide a framework where to easily plug-in custom modeling logic by extending the base classes. As of today, such classes include the Forecaster, the Reconstructor, and the ModelBasedAnomalyDetector. When extending base classes, a set of methods and decorators is automatically inherited, so that only the model-specific logic needs to be coded. For example, the base Forecaster class provides the apply(), evaluate() and cross_validate() methods, together with two decorators: @Forecaster.fit_method and @Forecaster.predict_method, to be used to decorate the custom fit() and predict() methods. All common functionalities, including the save() and load() methods as well as the data consistency checks, are thus automatically handled.
3 Illustrative examples
The following examples can be reproduced in a clean Timeseria installation, either in a Jupyter environment or with a standard Python script. In this case, plots must be generated as images and saved to file, adding the following extra arguments to each plot() call: image=True, save_to="plot.png".
3.1 Load some data
Load an air temperature dataset, sampled at about ten minutes intervals, as a TimeSeries object. Timestamps are assumed on UTC as not otherwise specified. The time series is then printed and plotted.
from timeseria import TEST_DATASETS_PATH as PATH
from timeseria.datastructures import TimeSeries
timeseries = TimeSeries.from_csv(PATH + "humitemp.csv")
print(timeseries)
timeseries.plot()
Time series of #14000 points at variable resolution (~615s), from point @
1546475294.0 (2019- 01-03 00:28:14 +00:00) to point @ 1555544819.0 (2019
-04-17 23:46:59 +00:00)
[INFO] Aggregating by "10" for improved plotting
Figure 2: The time series plotted with an automatic aggregation factor of ten. The lighter shadow underlying the line chart represents the dispersion of the original data.
3.2 Resample to hourly data
The time series loaded in the previous example is resampled at one hour sampling interval. Gaps are filled by linear interpolation and the data_loss index is added. Then, the time series in printed and plotted.
timeseries = timeseries.resample("1h")
print(timeseries)
timeseries.plot()
[INFO] Using auto-detected sampling interval: 615.0s
[INFO] Resampled 14000 DataTimePoints in 2519 DataTimePoints
Time series of #2519 points at 1h resolution, from point @ 1546477200.0 (2019
-01-03 01:00:00 +00:00) to point @ 1555542000.0 (2019-04-17 23:00:00 +00:00)
Figure 3: The resampled time series plotted together with the data loss index, which is rendered as a red area chart.
3.3 Three-days forecast
A forecasting model based on a LSTM neural network is first cross-validated, then fitted and applied with a forecasting horizon of 72 (hourly) steps, for a total of three days. Data marked as missing is excluded from both the training and the evaluation of the model. Results are then plotted.
from timeseria.models import LSTMForecaster
forecaster = LSTMForecaster(window=24, neurons=256,
    features=["values", "hours"])
print(forecaster.cross_validate(timeseries, rounds=3,
    evaluate_error_metrics=["RMSE", "MAPE"]))
forecaster.fit(timeseries, epochs=100, reproducible=True)
forecaster.apply(timeseries, steps=72).plot()
[INFO] Cross validation round 1/3: validate from 1546477200.0 (2019-01-03
01:00:00+00:00) to 1549497600.0 (2019-02-07 00:00:00+00:00), fit on the rest.
[INFO] Cross validation round 2/3: validate from 1549497600.0 (2019-02-07
00:00:00+00:00) to 1552518000.0 (2019-03-13 23:00:00+00:00), fit on the rest.
[INFO] Cross validation round 3/3: validate from 1552518000.0 (2019-03-13
23:00:00+00:00) to 1555538400.0 (2019-04-17 22:00:00+00:00), fit on the rest.
{’humidity[RH]_RMSE_avg’: 1.168,
 ’humidity[RH]_RMSE_stdev’: 0.207,
 ’humidity[RH]_MAPE_avg’: 0.0196,
 ’humidity[RH]_MAPE_stdev’: 0.004,
 ’temperature[C]_RMSE_avg’: 0.362,
 ’temperature[C]_RMSE_stdev’: 0.039,
 ’temperature[C]_MAPE_avg’: 0.011,
 ’temperature[C]_MAPE_stdev’: 0.002}
Figure 4: The time series plotted together with the three days (72 hourly steps) forecast, which is highlighted in yellow using a “forecast” data index.
3.4 Perform anomaly detection
This example fits and apply a simple anomaly detection model based on periodic averages. By default the anomaly detectors assumes to work in unsupervised mode, using some “sane defaults”. Data marked as missing is excluded from the anomaly detection process, in order to prevent false positives. The time series is then plotted together with the anomaly data index, which allow for visual inspection: the bigger anomaly is detected around the 25th of January, when there is an unexpected drop with respect to the “normal” behaviour.
from timeseria.models import PeriodicAverageAnomalyDetector
anomaly_detector = PeriodicAverageAnomalyDetector()
anomaly_detector.fit(timeseries, periodicity=24)
anomaly_detector.apply(timeseries).plot()
[INFO] Using a window of "24" for "humidity[RH]"
[INFO] Using a window of "24" for "temperature[C]"
[INFO] Predictive model(s) fitted, now evaluating...
[INFO] Computing actual vs predicted for "humidity[RH]"...
[INFO] Computing actual vs predicted for "temperature[C]"...
[INFO] Model(s) evaluated, now computing the error_metric distribution(s)...
[INFO] Anomaly detector fitted
Figure 5: The time series plotted together with the anomaly index: the closer the index is to the top of the plot (corresponding to the value “1”), the bigger the anomaly is.
3.5 Aggregate to daily data
This example aggregates the hourly time series in daily data, but first it changes the time zone to the right one for this data set. This is important for getting the daily aggregates right and properly handling the daylight saving time change in March. In the aggregation process, the minimum and maximum operations are also computed, besides the (default) average one. The time series is then plotted.
timeseries.change_tz("Europe/Rome")
timeseries.aggregate("1D", operations=["min","max","avg"]).plot()
[INFO] Using auto-detected sampling interval: 3600.0s
[INFO] Aggregated 2494 points in 103 slots
Figure 6: The time series aggregated in daily slots, together with the data loss index.
4 Impact
As of today, Timeseria is used for ongoing research in environmental science, astrophysics and economics, and in a few private companies in the water and energy management space, while being progressively rolled out.
We believe that Timeseria can impact many fields, and let researchers and professionals focus more on their core tasks and less on repetitive, low added-value ones, such as handling data consistency, time and calendar arithmetic, edge cases, model evaluation, inspection of the results, and more.
We also believe that a structured approach as proposed in this work can impact several aspects of time series analysis overall, as improving reproducibility, reducing false positives and negatives, preventing misleading conclusions, and in general to keep analysis process under control with the ultimate goal of fostering new research and applications.
5 Conclusions
We presented Timeseria, an object-oriented time series processing library implemented in Python. Our goal was to make it easier to manipulate time series data and to build statistical and machine learning models on top of it.
We tackled the various issues that arise in performing such tasks at their roots, defining a novel representation for time series data based on well defined logical units (objects), which we then used to build the library upon.
Thanks to this approach, we were able to handle most of the issues that arise in processing time series data by design, with particular attention to practical challenges and real-world scenarios.
Early adoption of Timeseria proved it to be a valid tool for everyday work, preventing both naive (yet hard to debug) and more substantial errors, while speeding up all the various activities.
Acknowledgements
Supported by the Italian Research Center on High Performance Computing, Big Data and Quantum Computing (ICSC), funded by the European Union - NextGenerationEU - and the Italian National Recovery and Resilience Plan (NRRP) - Mission 4, Component 2 - within the activities of the Spoke 3 (Astrophysics and Cosmos Observations)